# VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models

### v3

@@ Abstract
Most existing benchmarking approaches for evaluating the output quality of large language models (LLMs) rely on comparing LLM responses to predefined references. Such methods, based on static datasets, quickly become outdated as LLM capabilities and use cases evolve. In this work, we introduce VARCO Arena—a novel, cost-effective, and robust benchmarking approach that leverages a single-elimination tournament structure to minimize the number of required comparisons while eliminating the need for static references or costly human annotations. We validate our approach through two experiments: (i) a simulation study that examines its robustness under various conditions, and (ii) an empirical evaluation using publicly available benchmark prompts. In both experiments, VARCO Arena consistently outperforms current LLM benchmarking practices, achieving stronger correlations with human-established Elo ratings. Our results demonstrate that VARCO Arena not only produces reliable LLM rankings but also provides a scalable, adaptable solution for qualitative evaluation across diverse, customized use cases. We release our demo and code at [https://huggingface.co/spaces/NCSOFT/VARCO_Arena].
VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models
Seonil Son1, Ju-Min Oh1thanks: Current institution: Samsung AI Research, Heegon Jin1thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: deftson@ncsoft.com
@@ 1 Introduction
The versatility of Large Language Models (LLMs) stems from their generative capacity to address a wide array of tasks. The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios Ouyang et al. (2022); Köpf et al. (2024); Roziere et al. (2023). As versatility emerges as a core attribute of LLMs, the challenge of accurately gauging their skill becomes increasingly significant. In response to this challenge, numerous benchmarks evaluating LLM capabilities have emerged Hendrycks et al. (2020); Srivastava et al. (2023); Zhong et al. (2024). Many LLM benchmarks employ formats amenable to automated scoring. Examples include benchmarks for arithmetic problems Gao et al. (2022); Patel et al. (2021), multiple-choice questions Lin et al. (2022), and code execution Austin et al. (2021); Chen et al. (2021). While these benchmarks are valuable, they primarily assess problem-solving abilities. The need for LLM benchmarks extends beyond this, as the primary value of LLMs lies in the versatility of their generative behavior. Researchers have primarily relied on pairwise comparisons to evaluate and rank the LLMs based on human preference annotations. A representative example of this approach is Chatbot Arena Chiang et al. (2024), which computes Elo ratings based on a massive number of human votes. While Chatbot Arena benefited from reliable, open-ended prompts generated by a large user base, obtaining such a high volume of annotations from dedicated annotators remains challenging.
@@
Figure 1: VARCO Arena directly compares LLM response pairs in single-elimination tournament rather than comparing references. In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct head-to-head comparison is more intuitive and results in better separability.
No. Comp. ($\downarrow$) Judge Eval. Type
Chatbot Arena unknown human head-to-head
Current Practice $n_{\text{model}}\cdot|X|$ LLM reference -based
@@
VARCO Arena $(n_{\text{model}}-1)\cdot|X|$ LLM head-to-head
@@
Table 1: Comparison between Current Practice and Varco Arena. $|X|$ and $n_{\text{model}}$ represents size of benchmark dataset, and number of candidate LLMs to rank respectively. Human annotators are considered much more costly than LLM judge counterpart.
To address these challenges, several benchmark datasets for generation tasks have been developed and applied with an LLM judge Zheng et al. (2024) to quantify LLM capabilities. Using these benchmarks and their reference responses, LLMs are ranked via pairwise comparisons between their outputs and the reference responses provided by LLM judges. Representative examples include AlpacaEval Li et al. (2023), Arena-Hard-Auto Li et al. (2024), and MTBench Zheng et al. (2023). There are two major advantages to relying on reference responses: (1) The number of comparisons required for ranking scales linearly with the number of LLMs. (2) The reference response establishes a quality standard for evaluating LLM outputs.
However, we argue that using reference responses to mediate comparisons between LLM outputs is suboptimal and that head-to-head comparisons yield more accurate assessments. In this context, we propose VARCO Arena as a novel evaluation framework that directly compares LLM responses while requiring fewer comparisons. Even with fewer comparisons and without relying on reference responses, VARCO Arena achieves a stronger correlation with the human-established Elo rankings from Chatbot Arena.
VARCO Arena is designed to conduct a single-elimination tournament for each prompt across all participating LLM responses and then compute Elo ratings Elo and Sloan (1978) based on the outcomes of these matches. By organizing tournament-style matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs with fewer comparisons than those required by reference-based methods. These win rates are subsequently interpreted as Elo ratings, enabling us to generate a comprehensive ranking of candidate LLMs.
We validate the effectiveness of VARCO Arena against current LLM benchmarking practices through two experiments. First, we conduct a simulation study (Section 4.2) to evaluate the reliability and robustness of our tournament approach under various controlled factors, including the number of participating LLMs, the number of test prompts, and the accuracy of the judge model. Second, we empirically demonstrate that running VARCO Arena with several judge LLMs on public benchmark (Section 4.3) prompts yields a stronger correlation with human-established Elo rankings from Chatbot Arena than does the reference-based approach.
These experiments confirm that our benchmarking approach outperforms current practices in terms of ranking reliability while requiring the lesser number of comparisons (Table 1). Notably, the simulation results—which emulate outcomes under various conditions—underscore that the superiority of our tournament approach is not coincidental but rather reflects the inherent advantages of direct pairwise comparisons over mediated comparisons using reference responses.
In this work, we make the following contributions:
@@
    We introduce VARCO Arena, a novel reference-free benchmarking method for LLMs that leverages a single-elimination tournament structure to perform head-to-head comparisons of LLM responses across test prompts. This approach eliminates the dependency on static references and enhances benchmarking flexibility.
@@
    Through extensive simulation studies and empirical evaluations, we demonstrate that VARCO Arena not only yields more reliable LLM rankings compared to traditional reference-based methods but also achieves this with significantly fewer comparisons, thereby reducing evaluation costs.
@@
    We release our demo and code to facilitate the further research and reproducibility; our resources are available at [https://huggingface.co/spaces/NCSOFT/VARCO_Arena]
2 Preliminaries: Quantifying Generation Ability
Quantifying an LLM’s generation ability is crucial, but it presents several challenges. The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as the provided test prompts and the inherently subjective nature of human preferences. A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its real-world performance. Two popular measures for LLM generation ability are the win rate against reference responses and the Elo rating.
2.1 Win Rate in Comparison to the Reference
AlpacaEval Li et al. (2023) and Arena-Hard-Auto Li et al. (2024) are representative benchmark datasets that aim to quantify LLM response quality using reference responses. These benchmarks employ an LLM judge with a standardized prompt for automated evaluation of generation capability. Given the prompts, LLM responses, and reference responses, the judge LLM is tasked with determining whether the reference or the response is preferred. The LLM’s win rate across the test prompts in the benchmark is used as a measure of its generation proficiency.
2.2 Elo Rating
The Elo rating system, introduced by Elo and Sloan, has since become a popular method for quantifying performance levels in competitive sports and, more recently, for evaluating the generative abilities of LLMs. The primary purpose of the Elo rating system is to represent a participant’s skill level as a single scalar value, enabling the prediction of relative win rates between participants who have never directly competed. Assessing the superiority of one LLM over another in terms of generative capability shares several similarities with determining winners in competitive sports. Varying test prompts may yield different results, similar to how weather or other factors can influence the outcome of a sports match. While a superior team or LLM is not guaranteed to outperform its inferior counterpart in every instance, it tends to succeed more frequently. With Elo ratings, one can expect the relative chance of winning between a pair of players. The expected win rate of the player $i$ against $j$ is computed as follows:
@@ 2.2 Elo Rating
Computing accurate Elo ratings requires a sufficient number of matches to estimate relative win rates between pairs of participants. The resulting matrix of relative win rates is then used to estimate Elo ratings through logistic regression, similar to multivariate logistic regression with a sigmoid function. In Equation 1, $P(i>j)$ represents the expected win rate of participant $i$ over $j$, $R_{i}$ is the Elo rating of participant $i$, and $\Delta_{\text{ij}}$ is the Elo rating difference between participants $j$ and $i$ ($R_{j}-R_{i}$).
Chatbot Arena presents a leaderboard of LLMs, with Elo ratings computed from matches evaluated by a large user base. Users prompt a pair of LLMs and submit their judgments of which one responded better. Although Chatbot Arena relies on manual evaluation of matches, it offers an intuitive method for comparing LLMs using Elo ratings.
3 VARCO Arena
We propose VARCO Arena, a reference-free approach for benchmarking large language models (LLMs) via single-elimination tournaments. Instead of comparing model responses against fixed references, VARCO Arena directly compares outputs from different models, determining superiority through head-to-head matchups for each prompt in our benchmark dataset. Repeated tournaments across prompts yield reliable leaderboards that reflect the relative performance of each model.
We begin by motivating our approach over current reference-based evaluations (Section 3.1). Next, we detail how VARCO Arena performs tournaments and results in Elo ratings (Section 3.2 and Algorithm 1). Finally, we discuss the reason why aggregating the results from tournaments promises reliable ranking (Section 3.3).
3.1 Comparing to a Reference is not Always Helpful
Although reference texts are a standard way to evaluate and rank large language models (LLMs), they introduce potential failure modes. Beyond the fact that a single reference might not capture every dimension of correctness, relying solely on a reference can lead to unreliable rankings of LLMs.
Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs. If we choose to compare LLM responses directly to rank them using Elo ratings (Equation 1), all head-to-head comparisons are utilized. In contrast, reference-based evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
When the reference output ($Y_{i}$) for a prompt ($X_{i}$) successfully disambiguates the pair of LLM responses $M_{1}(X_{i})$ and $M_{2}(X_{i})$ (as in the first and second cases), comparison to the reference is effective for benchmarking. Otherwise, these comparisons do not help differentiate LLM performance. Consequently, the reference-based approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the reference.
3.2 Tournaments of LLMs over multiple prompts to Elo Ratings
Algorithm 1 Tournaments of LLMs over prompts
1:prompts $X=\{x_{1},~{}x_{2},~{}...,~{}x_{i}\}$, LLMs $M=\{m_{1},~{}m_{2},...,~{}m_{j}\}$, outputs $O_{i,j}$ = $m_{j}(x_{i})$
2:Ranked LLMs with Elo ratings
3:function Match($m_{1},m_{2},x$)
4:     return $m_{1}$ if $\texttt{IsBetter}(O_{x,1},O_{x,2})$
5:     else $m_{2}$
6:end function
7:function SingleElim($M$, $x$, res)
8:     if $|M|=2$ then
9:         res.append(Match($M$[0], $M$[1], $x$))
10:         return res[-1]
11:     end if
12:     mid $\leftarrow\lfloor|M|/2\rfloor$
13:     left $\leftarrow$ SingleElim($M$[:mid], $x$, res)
14:     right $\leftarrow$ SingleElim($M$[mid:], $x$, res)
15:     return SingleElim(left + right, $x$, res)
16:end function
17:function Tournaments2Ranks($X,M$)
@@ 3.2 Tournaments of LLMs over multiple prompts to Elo Ratings
19:     for $x_{i}\in X$ do
20:         SingleElim(Shuffled($M$), $x_{i}$, res)
21:     end for
22:     return ComputeElo(res)
23:end function
Figure 1 and Algorithm 1 illustrate how VARCO Arena benchmarks LLMs via a tournament approach. Here, $|X|$ denotes the number of prompts in the benchmark dataset. Each execution of VARCO Arena runs a tournament among participant LLMs for every prompt in the dataset.
The use of tournament structures for LLM benchmarking offers both benefits and challenges. A major advantage of a single-elimination tournament is efficiency. As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using references. However, single elimination tournament only identifies a champion, leaving the relative ordering of other participants unclear.
@@
    Aggregating match results allows for precise win rate estimation via Elo ratings, resulting in a well-aligned overall ranking.
@@
    More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt.
In Section 3.3, we further explain how aggregating multiple tournaments could yield an reliable ranking of LLMs. We also provide an analysis of the number of matches each LLM faces, offering a comprehensive view of the method’s efficiency and effectiveness.
3.3 Why Aggregating Multiple Tournaments Yields Reliable Ranks?
Our approach aggregates match outcomes from multiple tournaments to approximate the complete set of pairwise comparisons—akin to the comparisons made in merge sort. In a single-elimination tournament, every participant advances based solely on match outcomes, a process that mirrors the merging steps in merge sort. Notably, a single-elimination tournament executes only the comparisons strictly necessary for determining the winner, omitting many comparisons that merge sort would perform further.
We posit that the missing pairwise match-ups in any one tournament can be recovered by aggregating tournaments conducted over different prompts. This hypothesis relies on the assumption—central to our use of the Elo model—that match outcomes are independent of the prompt. Consequently, matches across different prompts are considered equivalent.
Considering only the initial comparisons, which are randomly sampled, the aggregate number of comparisons is $|X|\cdot n_{\text{model}}/2$. Since $n_{\text{model}}$ is typically on the order of tens and $|X|$ comprises at least hundreds, this number exceeds the total possible match-ups, $\binom{n_{\text{model}}}{2}$ [^1] Note that merge sort produces a set of match-ups with no duplicates.. Furthermore, as shown in Table 1, the remaining matches—totaling $|X|\cdot(n_{\text{model}}-1)$—either contribute additional merge sort structures or enhance the accuracy of estimating relative win rates among LLM participants.
Moreover, considering that each unique pair meets in at least $|X|/(2(n_{\text{model}}-1))$ matches [^2] A model participates in between $|X|$ and $\lceil\log_{2}n_{\text{model}}\rceil$ matches per tournament. Dividing this by the number of possible match-ups per model, $2(n_{\text{model}}-1)$, yields the expression above. across the benchmark, this frequency is sufficient for a fair estimation of their relative win rates.
In summary, aggregating tournaments not only reconstructs the full set of pairwise comparisons for a merge sort but also ensures that each pair of models faces one another often enough to yield accurate win rate estimations, leading to reliable Elo ratings. Based on these considerations, we propose that conducting tournaments over a benchmark prompts will yield a reliable ranking of LLMs.

### v4

# Arena-Lite: Efficient and Reliable Large Language Model Evaluation via Tournament-Based Direct Comparisons

@@ Abstract
As Large Language Models (LLMs) expand across domains, LLM judges have become essential for systems evaluation. Current benchmarks typically compare system outputs against baselines. This baseline-mediated approach, though convenient, yields lower reliability than direct comparison between systems. We propose Arena-Lite which integrates tournament structure on top of head-to-head comparison. The application of a tournament structure and direct comparison eliminates the need for baseline outputs, reduces the number of required comparisons, and allows higher reliability in system rankings. We conducted two experiments: (1) controlled stochastic modeling and (2) empirical validation with a real LLM judge. Those experiments collectively demonstrate that Arena-Lite consistently achieves higher reliability with fewer comparisons, even with smaller datasets or weaker judges. We release an easy-to-use web demonstration and code to foster adoption of Arena-Lite, streamlining model selection across research and industry communities. Arena-Lite demo and code are available on https://huggingface.co/spaces/NCSOFT/ArenaLite
Arena-Lite: Efficient and Reliable Large Language Model Evaluation via Tournament-Based Direct Comparisons
Seonil Son1thanks: Current institution: RLWRLD, Ju-Min Oh1thanks: Current institution: Samsung AI Research, Heegon Jin1thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: seonil.calendar@gmail.com
@@ 1 Introduction
Figure 1: Arena-Lite directly compares LLM response pairs over multiple single-elimination tournaments rather than comparing responses to baseline outputs. In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct head-to-head comparison is more intuitive and results in better separability.
Total no.
matches ($\downarrow$)
No. matches per
LLM participant ($\uparrow$)
[…]
@@
Arena-Lite (ours) $(n_{\text{model}}-1)\cdot|X|$ $[~|X|,~|X|*\lceil\log_{2}n_{\text{model}}\rceil~]$
@@
Table 1: Comparison between Current practice of benchmarking (comparing to baseline outputs) and Arena-Lite. $|X|$ and $n_{\text{model}}$ represents size of benchmark dataset, and number of candidate LLMs to rank respectively. Arena-Lite, always save |X| number of comparisons for benchmarking, while allows more matches per LLM participant thanks to head-to-head comparison.
LLMs excel in diverse tasks, from chatbots to code generation, due to their powerful generative capabilities Ouyang et al. (2022); Roziere et al. (2023). As their versatility grows, accurately evaluating their performance becomes critical. To address this, benchmarks like MMLU and BigBench have emerged to assess LLM capabilities across various domains Hendrycks et al. (2020); Srivastava et al. (2023). Many of these benchmarks, such as those for arithmetic or code execution (e.g., GSM-Hard, HumanEval Gao et al. (2022); Chen et al. (2021)), use automated scoring to evaluate problem-solving skills. However, their focus is not on quality of generated content or limited to the cases where the generated contents are automatically evaluated (e.g. programming), which are mostly not the case for variety of generation tasks. The Chatbot Arena, a leading platform for reliable human evaluation of LLMs, has set a standard by collecting extensive human annotations Chiang et al. (2024). Yet, its resource-intensive approach has prompted efforts to replicate its rankings using LLM judges as a cost-effective alternative Li et al. (2024, 2023). These methods, however, rely on baseline-mediated comparisons—comparing LLM outputs to a baseline encore model’s outputs—which sacrifice reliability.
Current benchmarks relying on encore models often rank LLMs by their win rate against baseline responses from an encore model. This approach has two advantages: it scales linearly with the number of LLMs and provides a consistent quality standard. However, we argue that comparing LLMs directly against each other is inherently more reliable than using baseline outputs, which can introduce noise coming from weak transitivity Xu et al. (2025) of human preferences on LLM responses. To address this, we propose Arena-Lite, a novel evaluation framework that uses direct, head-to-head comparisons organized in a tournament structure. By eliminating the need for baseline outputs, Arena-Lite reduces the number of comparisons required while achieving stronger alignment with human-established rankings, such as those from Chatbot Arena.
Arena-Lite conducts single-elimination tournaments among participating LLMs for each prompt. From the match results, we can compute Bradley-Terry preference ratings for the final ranking Bradley and Terry (1952). This results in a single scalar per model that captures relative performance between any model counterpart, enabling accurate and efficient ranking. We validate Arena-Lite through two experiments. The first experiment, stochastic modeling of LLM competition (Section 4.2) demonstrates that tournament-based direct comparison method outperforms baseline-mediated method under various conditions, including different numbers of LLM participants, dataset rows used, and judge accuracies. Second, our empirical experiment (Section 4.3) shows that Arena-Lite achieves higher correlation with Chatbot Arena’s rankings than standard approaches using baseline outputs (Table 1) attested toward number of LLM as judges. These results collectively highlight Arena-Lite’s ability to deliver reliable rankings with fewer comparisons, even with smaller datasets or weaker judges over various generation tasks.
Our contributions are threefold:
@@
    We introduce Arena-Lite, a tournament-based framework for direct LLM comparisons, offering greater reliability than baseline-mediated approaches.
@@
    We rigorously demonstrate, through both comprehensive modeling and empirical experiments, that Arena-Lite achieves more accurate LLM rankings while requiring fewer comparisons than prevalent practices, particularly those relying on common baseline model outputs.
@@
    We open-source a functional demo and the complete code for Arena-Lite (https://huggingface.co/spaces/NCSOFT/ArenaLite), enabling researchers and industry practitioners to easily host and utilize our framework for streamlined LLM evaluation.
2 Preliminaries: Quantifying Generative Performance
Quantifying the generative capabilities of LLMs is an inherently challenging task. The evaluation is complicated by the stochasticity of model outputs and the subjectivity of human judgments. To approximate real-world performance, a common methodology involves assessing model outputs across a diverse range of prompts. In this context, two metrics are widely utilized: the win rate, which measures the frequency of preference for a model’s output over a baseline, and the Bradley-Terry model, which is employed to infer a latent skill rating for each model based on pairwise comparisons.
2.1 Measuring Win rate over baseline outputs
Benchmarks like AlpacaEval and Arena-Hard-Auto assess LLM response quality by comparing it to baseline responses from a encore model Li et al. (2023, 2024). An LLM judge evaluates whether the candidate LLM’s response outperforms the baseline for a given prompt. The win rate—the proportion of prompts where the LLM’s response is preferred—serves as a measure of its generative ability. While this approach is straightforward and scalable, it introduces noise coming from mediated comparisons.
2.2 Bradley-Terry Model Preference for LLM Rating
The Bradley-Terry (BT) model Bradley and Terry (1952) is widely used to infer baseline-mediated rankings of LLMs from pairwise comparisons. Chatbot Arena adopts the BT model rather than the classical Elo system Elo and Sloan (1978), but both Elo and BT models are useful for expecting probability of match outcome based on a score difference, though they differ in update rules and statistical assumptions.
In the BT model, each LLM is assigned a latent score representing its procificency. Given LLMs $i$ and $j$ with scores $R_{i}$ and $R_{j}$, respectively, the probability that LLM $i$ is preferred over LLM $j$ is modeled as:
$\text{P}(i>j)=\frac{1}{1+10^{(R_{j}-R_{i})/400}}.$ (1)
This formulation closely resembles the Elo win-probability function, reinforcing the intuitive connection between the two.
Chatbot Arena uses this BT-based formulation to rank LLMs by aggregating human preferences collected through pairwise matchups Chiang et al. (2024). Users are shown responses from two anonymized models to the same prompt and asked to select which response they prefer. The accumulated judgments are then used to fit BT scores, producing a leaderboard that reflects relative model performance.
While this approach requires a substantial number of human evaluations to ensure reliability, it captures nuanced quality differences between models more effectively than purely automatic benchmarks. Arena-Lite, introduced in the next section, builds on the same BT modeling framework but seeks to reduce the number of required comparisons by using tournament-structured match-making.
3 Arena-Lite
To address the high annotation cost of Chatbot Arena while preserving evaluation reliability, we propose Arena-Lite. Arena-Lite introduces a tournament-based approach for efficient and reliable LLM evaluation using a single-elimination structure. Unlike baseline-mediated evaluations that compare model outputs to a baseline, Arena-Lite directly compares outputs from different models through head-to-head matchups for each prompt in benchmark datasets. Repeated tournaments across the dataset produce consistent leaderboards reflecting models’ fundamental performance.
We first discuss limitations of baseline-mediated evaluations (Section 3.1). Next, we describe how Arena-Lite conducts tournaments to generate ratings (Section 3.2, Algorithm 1). Finally, we highlight similarities between the single-elimination structure and merge sort, explaining why aggregated tournaments yield reliable LLM rankings (Section 3.3).
3.1 Comparing to Baseline outputs is not Always Helpful
Although baseline outputs are a standard way to evaluate and rank LLMs, they introduce potential failure modes. Beyond the fact that a single baseline output might not capture every dimension of appropriate answers, relying solely on a baseline output can lead to unreliable rankings of LLMs.
Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs. If we choose to compare LLM responses directly to rank them using BT preference (Equation 1), all head-to-head comparisons are utilized. In contrast, baseline-mediated evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
When the baseline output ($Y_{i}$) for a prompt ($X_{i}$) successfully disambiguates the pair of LLM responses $M_{1}(X_{i})$ and $M_{2}(X_{i})$ (as in the first and second cases), comparison to the baseline is effective for benchmarking. Otherwise, these comparisons do not help differentiate LLM performance. Consequently, the baseline-mediated approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the baseline.
3.2 Tournaments of LLMs over multiple prompts to preference ratings
Algorithm 1 Tournament-Based Model Evaluation
1:Models $M=\{m_{1},m_{2},\ldots,m_{n}\}$, Prompts $X=\{x_{1},x_{2},\ldots,x_{k}\}$
2:Bradley-Terry preference ratings
3:Initialize $R\leftarrow\emptyset$
4:for each $x_{j}\in X$ do
@@ 3.2 Tournaments of LLMs over multiple prompts to preference ratings
7:  $M^{\prime}\leftarrow M\cup\{\texttt{None}\}^{\texttt{n\_byes}}$
8:  Randomly shuffle $M^{\prime}$
[…]
10:end for
[…]
13:  if $|M^{\prime}|=2$ then
[…]
15:   if $\texttt{None}\notin M^{\prime}$ then
16:     Add result to $R$
17:   end if
[…]
20:   end if
21:  else
22:   $\texttt{mid}\leftarrow\lfloor|M^{\prime}|/2\rfloor$
[…]
25:  end if
26:end function
[…]
28:  if $m_{i}=\texttt{None}$ and $m_{j}=\texttt{None}$ then
29:   return None
30:  else if $m_{i}=\texttt{None}$ or $m_{j}=\texttt{None}$ then
31:   return $m_{i}$ if $m_{j}=\texttt{None}$ else $m_{j}$
32:  else
33:   $O_{i}\leftarrow m_{i}(x)$, $O_{j}\leftarrow m_{j}(x)$
34:   if $O_{j}>O_{i}$ then return $(m_{j},m_{i})$
35:   elsereturn $(m_{i},m_{j})$
36$\triangleright$ returns (winner, loser) tuple
37:   end if
38:  end if
39:end function
Figure 1 and Algorithm 1 illustrate how Arena-Lite benchmarks LLMs via a tournament approach. Here, $|X|$ denotes the number of prompts in the benchmark dataset. Running Arena-Lite hosts tournaments among participant LLMs for every prompt in the dataset.
The use of tournament structures for LLM benchmarking offers both benefits and challenges. A major advantage of a single-elimination tournament is efficiency. As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using baseline outputs. However, single elimination tournament only identifies a champion, leaving the relative ordering of other participants unclear.
@@
    Aggregating match results allows for precise win rate estimation via BT preference, resulting in a well-aligned overall ranking.
@@
    More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt (Table 1).
In Section 3.3, we further explain how aggregating multiple tournaments could yield reliable ranking of LLMs. We also provide further analysis on number of comparisons performed over tournaments of Arena-Lite comparing to merge sort, offering a comprehensive view of the method’s efficiency and effectiveness.
3.3 Why Aggregating Multiple Tournaments Yields Reliable Rankings
[…]
To achieve reliable rankings of LLMs, our approach aggregates match outcomes from multiple tournaments, lesser than a full grid comparisons, but effectively approximating the complete set of pairwise comparisons required for merge sort. We outline the rationale in four key points:
Merge Sort Baseline A single-elimination tournament mirrors the merging steps of merge sort, which requires $\mathcal{O}(n\log n)$ comparisons with no duplicate match-ups to rank $n$ models. However, a single tournament omits many comparisons, covering only the minimal match-ups needed to determine a winner.
Recovering Comparisons via Aggregation By aggregating tournaments over diverse prompts, we can recover missed pairwise match-ups that should have occurred. Assuming match outcomes are prompt-independent (as BT model assumes), matches across prompts are considered equivalent. With $|X|$ prompts (typically hundreds to thousands) and $n_{\text{model}}$ models (tens), only random initial match-ups totaling $|X|\cdot\lfloor\frac{n_{\text{model}}}{2}\rfloor$. This exceeds the $\binom{n_{\text{model}}}{2}$ possible combination, ensuring broad coverage.
Sufficiency of Comparisons The aggregated match-ups not only cover the necessary comparisons but also surpass the $\mathcal{O}(n\log n)$ requirement of merge sort. Moreover, each unique model pair competes in average $|X|/(n_{\text{model}}-1)$ to $\frac{|X|\log_{2}n_{\text{model}}}{n_{\text{model}}-1}$ matches [^1] These estimates are computed by number of matches a model undergoes which is $[~|X|,~|X|*\lceil\log_{2}n_{\text{model}}\rceil~]$, devided by number of possible opponents for a model, $n_{\text{model}}-1$ across the benchmark, a frequency mostly sufficient for accurate win rate estimation.
Refinement for Reliability The remaining matches, totaling $|X|\cdot(n_{\text{model}}-1)$, further refine the ranking by enhancing win rate estimates, especially among top-performing models, reducing noise and ensuring robustness akin to Arena-Lite’s sampling strategy.
In summary, aggregating multiple tournaments reconstructs the full set of comparisons needed for a merge sort-like ranking while providing enough repeated match-ups to ensure accurate win rate estimations. This dual mechanism yields reliable and robust LLM rankings across the benchmark.